OpenStudy (anonymous):

In her last four games of golf, Maria has had scores of 76, 73, 70, and 67. If the pattern continues, what will Maria's score be in her next game?

OpenStudy (turingtest):

76-3=73 73-3=70 70-3=67 ... continue...

OpenStudy (anonymous):

Why would u subtract it by 3? I thought u would subtract it by 4?..

OpenStudy (anonymous):

:c Can I please just get an answer as to why?

OpenStudy (turingtest):

Because each successive number is 3 less than the one before it... to continue the pattern you would continue to subtract by 3 to find the next number.

OpenStudy (anonymous):

This is arithmetic progression with common difference 3.

OpenStudy (anonymous):

a*

OpenStudy (anonymous):

I did I just wanted to know why o.o ty

OpenStudy (anonymous):

Yes that's the reason, read my post again and again ...

OpenStudy (anonymous):

Thank you ^-^ Trust me I have too. For the unit I am currently working on. (online course for what I forgot) D: